# CHEMOUT: CHEMical complexity in star-forming regions of the OUTer Galaxy

### v2

@@ Appendix B Millimeter continuum maps and H2 column densities
where $\nu_{2}$ and $\nu_{1}$, are 150 and 260 GHz (i.e., the frequencies corresponding to 2.0 and 1.2 mm), and $T_{\text{d}}$ is the assumed dust temperature (see Sect. 3.2). In Fig. 8, the galactocentric trends of H2 are shown. The dataset estimated with a non-constant gas-to-dust ratio $\gamma=\gamma(R_{\text{GC}})$ (giannetti2017galactocentric) exhibits a positive trend with increasing values of $R_{\text{GC}}$, and a weak, positive correlation ($\rho=0.19$), which is due to the increasing $\gamma$ (Sect. 3.2).
@@
Figure 8: Gradients of column densities of H2, $N_{\text{H}_{2}}$, as a function of galactocentric radius ($R_{\text{GC}}$). The light blue dataset represents the column densities estimated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue dataset shows the column densities calculated with a gas-to-dust ratio function of $R_{\text{GC}}$ from giannetti2017galactocentric. Both datasets have been fitted with a linear regression. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. In the upper-right side of the plot, the Pearson correlation coefficient, $\rho$, is shown (only for the dark blue dataset).
@@ Appendix C Line-fitting results
In this appendix, we present the results of the line fit from MADCUBA. Tables 7 and 8 present the best-fit parameters from MADCUBA for HCN, HCO+, c-C3H2, H13CO+, HCO, and SO. The fit results for H2CO and CH3OH have already been presented in fontani2022chemoutII. The upper-limits values are also estimated for the undetected transitions. For the (self-) absorbed lines, no best-fit parameters are given.
In Figs. 9, 10, and 11, the observed spectra of the whole sample are shown with the Gaussian fits performed with MADCUBA for HCN, H13CO+, and SO. The plots showing the spectra of the remaining species are shown in fontani2022chemout (c-C3H2, HCO+) and in fontani2022chemoutII (HCO, H2CO, CH3OH).
@@ Appendix E Galactocentric gradients of H13CN
Figure 13: Galactocentric gradients of fractional abundances w.r.t. H2, $X_{\text{mol}}$, of H13CN. The light blue data represent the abundances calculated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue data illustrate those estimated using a non-constant gas-to-dust ratio ($\gamma=\gamma(R_{\text{GC}})$ from giannetti2017galactocentric). For both datasets, the linear regression results are shown as the light blue and dark blue lines, respectively. The $1\sigma$ error bars over the slope of the gradients are displayed for each molecular fit. The upper limit values are represented with triangles. The gradients of the elemental abundances of carbon (13C) and nitrogen (N), as reported by mendez2022gradients, are plotted as dashed lines. The product of the parent elements of the species is represented by a black dash–dot line. All the elemental trends are plotted shifted to the start of the molecular gradients estimated with $\gamma(R_{\text{GC}})$ for reference to allow comparison of their slopes with the linear fit of the molecular abundances. In the upper-right side of each plot, the Pearson correlation coefficient, $\rho$, is shown (estimated only for the abundances estimated using the non-constant gas-to-dust ratio).

### v3

@@ Appendix B Millimeter continuum maps and H2 column densities
where $\nu_{2}$ and $\nu_{1}$, are 150 and 260 GHz (i.e., the frequencies corresponding to 2.0 and 1.2 mm), and $T_{\text{d}}$ is the assumed dust temperature (see Sect. 3.2). In Fig. 8, the galactocentric trends of H2 are shown. The dataset estimated with a non-constant gas-to-dust ratio $\gamma=\gamma(R_{\text{GC}})$ (Giannetti et al. 2017) exhibits a positive trend with increasing values of $R_{\text{GC}}$, and a weak, positive correlation ($\rho=0.19$), which is due to the increasing $\gamma$ (Sect. 3.2).
@@
Figure 8: Gradients of column densities of H2, $N_{\text{H}_{2}}$, as a function of galactocentric radius ($R_{\text{GC}}$). The light blue dataset represents the column densities estimated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue dataset shows the column densities calculated with a gas-to-dust ratio function of $R_{\text{GC}}$ from Giannetti et al. (2017). Both datasets have been fitted with a linear regression. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. In the upper-right side of the plot, the Pearson correlation coefficient, $\rho$, is shown (only for the dark blue dataset).
@@ Appendix C Line-fitting results
In this appendix, we present the results of the line fit from MADCUBA. Tables 7 and 8 present the best-fit parameters from MADCUBA for HCN, HCO+, c-C3H2, H13CO+, HCO, and SO. The fit results for H2CO and CH3OH have already been presented in Fontani et al. (2022b). The upper-limits values are also estimated for the undetected transitions. For the (self-) absorbed lines, no best-fit parameters are given.
In Figs. 9, 10, and 11, the observed spectra of the whole sample are shown with the Gaussian fits performed with MADCUBA for HCN, H13CO+, and SO. The plots showing the spectra of the remaining species are shown in Fontani et al. (2022a) (c-C3H2, HCO+) and in Fontani et al. (2022b) (HCO, H2CO, CH3OH).
@@ Appendix E Galactocentric gradients of H13CN
Figure 13: Galactocentric gradients of fractional abundances w.r.t. H2, $X_{\text{mol}}$, of H13CN. The light blue data represent the abundances calculated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue data illustrate those estimated using a non-constant gas-to-dust ratio ($\gamma=\gamma(R_{\text{GC}})$ from Giannetti et al. 2017). For both datasets, the linear regression results are shown as the light blue and dark blue lines, respectively. The $1\sigma$ error bars over the slope of the gradients are displayed for each molecular fit. The upper limit values are represented with triangles. The gradients of the elemental abundances of carbon (13C) and nitrogen (N), as reported by Méndez-Delgado et al. (2022), are plotted as dashed lines. The product of the parent elements of the species is represented by a black dash–dot line. All the elemental trends are plotted shifted to the start of the molecular gradients estimated with $\gamma(R_{\text{GC}})$ for reference to allow comparison of their slopes with the linear fit of the molecular abundances. In the upper-right side of each plot, the Pearson correlation coefficient, $\rho$, is shown (estimated only for the abundances estimated using the non-constant gas-to-dust ratio).